OpenStudy (anonymous):

solve symbolically 3 sq x-3+27=51

OpenStudy (anonymous):

(x-3) is in a radical?

myininaya (myininaya):

I think thats what he means

OpenStudy (anonymous):

it is the radical

OpenStudy (anonymous):

\[3\sqrt{x-3}+27=51\]

OpenStudy (anonymous):

move the 27 over, then divide by 3 on both sides

OpenStudy (anonymous):

51+27 then divide by 3

OpenStudy (anonymous):

no, subtract 27 from both sides

OpenStudy (anonymous):

u get 24, then divide by 3 to get rid of the number infront of the radical.. so u get 8

OpenStudy (anonymous):

yeah that is what i have

OpenStudy (anonymous):

so what then do I put it \[3\sqrt{8-3?}\]

OpenStudy (anonymous):

27 cancels on left side

OpenStudy (anonymous):

\[3\sqrt{x-3}=24\]

OpenStudy (anonymous):

\[\sqrt{x-3}=24/3\]

OpenStudy (anonymous):

\[\sqrt{x-3}=8\]

OpenStudy (anonymous):

that is the answer? symbolically

OpenStudy (anonymous):

r u trying to find x? is that what u mean by symbolically

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

all it says is solve the equation symbolically

OpenStudy (anonymous):

\[(\sqrt{x-3})^2= (8)^2\]

OpenStudy (anonymous):

x-3=64, add 3 to the other side, nd u get 67

OpenStudy (anonymous):

okay I got that i see where you get it

OpenStudy (anonymous):

that is what x is equal to, but i dont know what it means by solve symbolically..i hope it helped a bit

OpenStudy (anonymous):

thank you I will see when I do my test and submit it